OpenStudy (anonymous):

Which equation is in slope-intercept form & represents a line with slope 3 through the point (9, -4)? A) y = 3x + 21 B) -3x + y = -31 C) y = 3x - 31 D) -3x + y = 21

OpenStudy (anonymous):

I know that B & D aren't possibilities.

OpenStudy (anonymous):

@tanner23456

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@mathstudent55

OpenStudy (tanner23456):

so slope or m = 3 I usually like to plot it on the graph and find where the line crosses the y-axis to find the y-intercept but you can also plug it into the point slope form y - y1 = m(x-x1) y - (-4) = 3(x-9) it becomes y + 4 = 3x - 27 subtract the 4 y = 3x - 31 C looks like the correct answer

OpenStudy (anonymous):

I thought the answer was A but thank you for showing me how you worked that out.